# Gaussian Process Learning of Nonlinear Dynamics

### v1

@@
Abstract: One of the pivotal tasks in scientific machine learning is to represent underlying dynamical systems from time series data. Many methods for such dynamics learning explicitly require the derivatives of state data, which are not directly available and can be approximated conventionally by finite differences. However, the discrete approximations of time derivatives may result in a poor estimation when state data are scarce and/or corrupted by noise, thus compromising the predictiveness of the learned dynamical models. To overcome this technical hurdle, we propose a new method that learns nonlinear dynamics through a Bayesian inference of characterizing model parameters. This method leverages a Gaussian process representation of states, and constructs a likelihood function using the correlation between state data and their derivatives, yet prevents explicit evaluations of time derivatives. Through a Bayesian scheme, a probabilistic estimate of the model parameters is given by the posterior distribution, and thus a quantification is facilitated for uncertainties from noisy state data and the learning process. Specifically, we will discuss the applicability of the proposed method to two typical scenarios for dynamical systems: parameter identification and estimation with an affine structure of the system, and nonlinear parametric approximation without prior knowledge.
@@ 1 Introduction
Ideally, physical consistency should be preserved in the learning process through a proper system parametrization, so that non-physical behaviors in the data are filtered out and reliability in long-term predictions is achieved. One popular choice is the sparse identification of nonlinear dynamics (SINDy) [7], which defines $f_{i}(\cdot;{\bf\it\theta}_{i})$ as a linear combination of candidate functions and uses sparse regression to determine active terms among them. The effectiveness of identification in this method relies on the inclusiveness of candidates and the assumption of sparsity. Relevantly, data-driven operator inference [25, 27, 17] focuses on reduced-order dynamics learning, and attempts to recover a low-dimensional system (1) with a polynomial structure $f_{i}(\cdot;{\bf\it\theta}_{i})$ inspired by known governing equations. Formulated as (Tikhonov-regularized) least squares [22], such a non-intrusive method seeks to find estimators of ${\bf\it\theta}_{i}$ that best match the training data in a minimum-residual sense, and hence avoids entailing access to source code. This is especially advantageous for applications with readily-executed solvers that cannot be easily modified. Instead of pinpointing the exact structure of a dynamical system, alternatively, $f_{i}(\cdot;{\bf\it\theta}_{i})$ can be considered as a parametric approximation of the system description. For example, neural ordinary differential equations [9] adopt a multi-layer perception network as $f_{i}(\cdot;{\bf\it\theta}_{i})$, while the Runge-Kutta neural network method [38] constructs parametric surrogate models for the discrete-time integrator.
@@
However, the predictive performance of these dynamics learning techniques may be compromised when training data are scarce and/or corrupted by noise. In particular for the minimum-residual of the continuous-time formulation (1), there is often no direct access to the required time derivatives, and they have to be approximated by numerical differentiation (e.g., finite differences). The data noise and/or sparsity would lead to a poor estimation of time derivatives and then curtail the accuracy of parameter inference. Efforts have been made to avoid derivative evaluations by considering state data misfit directly instead of minimizing residuals, such as incorporating time integration into loss functions [9] or as constraints [33]. However, these settings typically demand more sophisticated numerical treatments, which are not always convenient or effective for complex dynamical behaviors. Meanwhile, particularly for the cases with imperfect datasets, it is crucial to facilitate the learning process with uncertainty quantification to enable a model validation in terms of accuracy and robustness. A straightforward strategy is ensemble learning that leverages bootstrap sampling of time-series data to aggregate statistical estimates from multiple models [14], but the many times of evaluations may escalate computational costs. An efficient and rigorous alternative is Bayesian inference [6] for system parameters ${\bf\it\theta}_{i}$, deemed capable of taking data and modeling uncertainties into systematical consideration [31, 37, 35, 19, 17]. In this work, we aim to address the aforementioned two challenges – robustness with scarce/noisy data and capability of uncertainty quantification – through a novel Bayesian inference method for learning dynamical systems (1). The primary novelty lies in integrating Gaussian process modeling into Bayesian inference to account the approximation and uncertainty estimation of time derivatives, both accurately and efficiently.
@@
Gaussian process emulation is a probabilistic approach to supervised learning [29], and its variations have been applied to a wide range of computational tasks in science and engineering, see [4, 21, 13, 20, 8, 12, 36, 5, 11] for a few examples. More relevant to this work is the fact that Gaussian processes are also capable of embedding differential equation constraints into kernel structures. Specifically, as Gaussian properties can be preserved through linear operations, the latent force models [39] for Gaussian processes have paved the way for describing the correlations between solution states/fields and their latent ‘forces’ through linear differential operators. This idea has been used for solving forward and inverse problems governed by linear partial differential equations [28, 26], and extended to nonlinear equations [10, 23] with error analysis [1], as well as to nonlinear operator learning [2]. Relevant techniques have been explored in various contexts in scientific computing, such as stochastic partial differential equations [30], statistical finite element analysis [16], port-Hamiltonian systems [3], and conservation laws [18].
@@
In this work, a vector-valued Gaussian process is constructed to represent the statistical correlation between the states and their time derivatives. Combined with the system parametrization $f_{i}(\cdot;{\bf\it\theta}_{i})$, this formulates a likelihood function constrained by the differential equations (1), in addition to a prior distribution of ${\bf\it\theta}_{i}$ needed for the Bayesian inference of these parameters. Thereafter, a probabilistic estimate of ${\bf\it\theta}_{i}$ is given by the posterior distribution through the Bayes’ rule. By sampling over this distribution, uncertainties are consequently propagated into the predictions produced by the learned dynamics system, enabling a predictive uncertainty quantification. Importantly, the Gaussian process representation also provides a natural smoother by kernel regression, which counters the compromise of time derivative accuracy with noisy/scarce state data. We will demonstrate the proposed method in two scenarios: one with a given affine structure of the dynamical system, and the other for a general, nonlinear parametric approximation of the system without prior knowledge.
@@ 2 Dynamics learning with Gaussian processes
$p({\bf\it\theta}_{i})\propto\exp\left(-\frac{1}{2}\|{\bf\it\theta}_{i}\|_{{\bf\it\Lambda}_{i}}^{2}\right)\,,$ (2)
@@ 2.1 Likelihood definition with a Gaussian process
Here we assume that each state variable $x_{i}(t)$ ($1\leq i\leq N$) follows a Gaussian process:
@@
in which $\kappa_{i}$ is a given smooth kernel function representing the covariance over the Gaussian process. In particular, we use the squared exponential kernel in this work. Thus, the time-derivative of $x_{i}(t)$ also follows a Gaussian process, and the correlation between the state $x_{i}$ and its time-derivative $\dot{x}_{i}$ is given by
@@
For notation, we collect the observed data of $x_{i}(t)$ at the time-instances $\mathcal{T}$ (i.e., $\{x_{i}(t_{k})\}_{k=0}^{K-1}$) in a vector $\mathbf{u}_{i}\in\mathbb{R}^{K}$, let $\mathbf{d}_{i}\in\mathbb{R}^{K}$ denote the time derivatives of $x_{i}$ over $\mathcal{T}$, and let $\mathbf{U}=[\mathbf{u}_{1}~{}\cdots~{}\mathbf{u}_{N}]^{\mathsf{T}}=[{\bf\it x}(t_{0})~{}\cdots~{}{\bf\it x}(t_{K-1})]\in\mathbb{R}^{N\times K}$ collect the full-state data of ${\bf\it x}(t)$ over $\mathcal{T}$. As a Gaussian process assumption on $x_{i}(t)$ has been adopted, $\mathbf{d}_{i}$ combined with $\mathbf{u}_{i}$ should follow a joint normal distribution, and (4) thus gives
$\begin{split}p\left(\mathbf{d}_{i}=f_{i}(\mathbf{U};{\bf\it\theta}_{i}),\mathbf{u}_{i}|~{}{\bf\it\theta}_{i}\right)&\propto\exp\left(-\frac{1}{2}\begin{bmatrix}f_{i}(\mathbf{U};{\bf\it\theta}_{i})\\ \mathbf{u}_{i}\end{bmatrix}^{\mathsf{T}}\begin{bmatrix}\mathbf{K}^{dd}_{i}&\mathbf{K}^{du}_{i}\\ \mathbf{K}^{ud}_{i}&\mathbf{K}^{uu}_{i}\end{bmatrix}^{\text{-1}}\begin{bmatrix}f_{i}(\mathbf{U};{\bf\it\theta}_{i})\\ \mathbf{u}_{i}\end{bmatrix}\right)\,,\\ \mathbf{K}_{i}^{dd}&=\partial_{t}\partial_{t^{\prime}}\kappa_{i}(\mathcal{T},\mathcal{T})+\chi^{d}_{i}\mathbf{I}_{K}\in\mathbb{R}^{K\times K}\,,\\ \mathbf{K}_{i}^{uu}&=\kappa_{i}(\mathcal{T},\mathcal{T})+\chi^{u}_{i}\mathbf{I}_{K}\in\mathbb{R}^{K\times K}\,,\\ \mathbf{K}_{i}^{du}&=\partial_{t}\kappa_{i}(\mathcal{T},\mathcal{T})=(\mathbf{K}_{i}^{ud})^{\mathsf{T}}\in\mathbb{R}^{K\times K}\,.\end{split}$ (5)
Considering the possibility that the data $\mathbf{U}$ may be corrupted by noise, white noise terms with variance values $\chi^{u}_{i}>0$ and $\chi^{d}_{i}>0$ are added to $\mathbf{u}_{i}$ and $\mathbf{d}_{i}$, respectively. This equation defines a likelihood, not only on the direct observations $\mathbf{u}_{i}$, but also on the alignment between $\dot{x}_{i}(\mathcal{T})=\mathbf{d}_{i}$ and $f_{i}(\mathbf{U};{\bf\it\theta}_{i})$, i.e., the satisfaction of the dynamical system formulation (1). Therefore, in addition to the data-driven nature, such a likelihood definition guarantees a physics-based interpretation.
Remark 1: In this work, we consider both $\mathbf{u}_{i}$ and $\mathbf{d}_{i}$ over the whole data coverage $\mathcal{T}$, but this is not a requirement, i.e., $\mathbf{u}_{i}$ and $\mathbf{d}_{i}$ can correspond to two different subsets of time instances in $\mathcal{T}$.
@@ 2.2 Posterior distribution for Bayesian inference
Note that
@@
Therefore, the posterior distribution can be written explicitly as
$\begin{split}&~{}~{}p\left({\bf\it\theta}_{i}|~{}\mathbf{d}_{i}=f_{i}(\mathbf{U};{\bf\it\theta}_{i}),\mathbf{u}_{i}\right)\propto p\left(\mathbf{d}_{i}=f_{i}(\mathbf{U};{\bf\it\theta}_{i}),\mathbf{u}_{i}|~{}{\bf\it\theta}_{i}\right)p({\bf\it\theta}_{i})\\ \propto&~{}\exp\left(-\frac{1}{2}\left(f_{i}(\mathbf{U};{\bf\it\theta}_{i})^{\mathsf{T}}\mathbf{R}_{i}^{dd}f_{i}(\mathbf{U};{\bf\it\theta}_{i})+2f_{i}(\mathbf{U};{\bf\it\theta}_{i})^{\mathsf{T}}\mathbf{R}_{i}^{du}\mathbf{u}_{i}+\|{\bf\it\theta}_{i}\|_{{\bf\it\Lambda}_{i}}^{2}\right)\right)\,.\end{split}$ (8)
Furthermore, we define an estimate of the time derivatives over $\mathcal{T}$ (i.e., $\dot{x}_{i}(\mathcal{T})$) using a Gaussian process regression with the state observations $(\mathcal{T},\mathbf{u}_{i})$ at the same instances, given as
@@
Considering the fact that $\mathbf{R}_{i}^{dd}\hat{\mathbf{d}}_{i}=-\mathbf{R}_{i}^{du}\mathbf{u}_{i}$, the posterior expression in (8) is hence rearranged into
$p\left({\bf\it\theta}_{i}|~{}\mathbf{d}_{i}=f_{i}(\mathbf{U};{\bf\it\theta}_{i}),\mathbf{u}_{i}\right)\propto\exp\left(-\frac{1}{2}\left(\left\|f_{i}(\mathbf{U};{\bf\it\theta}_{i})-\hat{\mathbf{d}}_{i}\right\|_{\mathbf{R}_{i}^{dd}}^{2}+\|{\bf\it\theta}_{i}\|_{{\bf\it\Lambda}_{i}}^{2}\right)\right)\,,$ (10)
@@
$({\bf\it\theta}_{i})_{\text{MAP}}=\arg\min_{{\bf\it\theta}_{i}\in\mathbb{R}^{p_{i}}}\left\{\left\|f_{i}(\mathbf{U};{\bf\it\theta}_{i})-\hat{\mathbf{d}}_{i}\right\|_{\mathbf{R}_{i}^{dd}}^{2}+\|{\bf\it\theta}_{i}\|_{{\bf\it\Lambda}_{i}}^{2}\right\}\,.$ (11)
This is a regularized, generalized least squares estimator of ${\bf\it\theta}_{i}$ weighted by $\mathbf{R}_{i}^{dd}=(\mathbf{K}^{dd}_{i}-\mathbf{K}^{du}_{i}(\mathbf{K}^{uu}_{i})^{-1}\mathbf{K}^{ud}_{i})^{-1}$, the inverse of which is the posterior covariance matrix of $\dot{x}_{i}(\mathcal{T})$ through Gaussian process regression.
@@
$p\left({\bf\it\theta}_{i}|~{}\mathbf{d}_{i}=f_{i}(\mathbf{U};{\bf\it\theta}_{i}),\mathbf{u}_{i}\right)\propto\exp\left(-\frac{1}{2}\left(\sum_{\texttt{all ICs}}\left\{\left\|f_{i}(\mathbf{U};{\bf\it\theta}_{i})-\hat{\mathbf{d}}_{i}\right\|_{\mathbf{R}_{i}^{dd}}^{2}\right\}_{\texttt{each IC}}+\|{\bf\it\theta}_{i}\|_{{\bf\it\Lambda}_{i}}^{2}\right)\right)\,,$ (12)
where the loss terms $\|f_{i}(\mathbf{U};{\bf\it\theta}_{i})-\hat{\mathbf{d}}_{i}\|_{\mathbf{R}_{i}^{dd}}^{2}$ are summed over all the initial conditions (ICs) included in the dataset. Note in this case that the trajectory with each initial condition is approximated by an individual GP, leading correspondingly to an individual $\mathbf{R}_{i}^{dd}$ as well. In particular, when data with a single initial condition are scarce, one can try to improve the parameter inference by adding data with other initial conditions. The estimated parameter values should work for the predictions for new initial conditions outside the coverage of training data.
@@
$\begin{split}&p\left({\bf\it\theta}_{m}\cup{\bf\it\theta}_{n}|~{}\mathbf{d}_{m}=f_{m}(\mathbf{U};{\bf\it\theta}_{m}),\mathbf{d}_{n}=f_{n}(\mathbf{U};{\bf\it\theta}_{n}),\mathbf{u}_{m},\mathbf{u}_{n}\right)\\ \propto~{}&\exp\left(-\frac{1}{2}\left(\left\|f_{m}(\mathbf{U};{\bf\it\theta}_{m})-\hat{\mathbf{d}}_{m}\right\|_{\mathbf{R}_{m}^{dd}}^{2}+\left\|f_{n}(\mathbf{U};{\bf\it\theta}_{n})-\hat{\mathbf{d}}_{n}\right\|_{\mathbf{R}_{n}^{dd}}^{2}+\|{\bf\it\theta}_{i}\|_{{\bf\it\Lambda}_{m}}^{2}+\|{\bf\it\theta}_{i}\|_{{\bf\it\Lambda}_{n}}^{2}\right)\right)\,.\end{split}$ (13)

### v2

@@
Abstract: One of the pivotal tasks in scientific machine learning is to represent underlying dynamical systems from time series data. Many methods for such dynamics learning explicitly require the derivatives of state data, which are not directly available and can be approximated conventionally by finite differences. However, the discrete approximations of time derivatives may result in poor estimations when state data are scarce and/or corrupted by noise, thus compromising the predictiveness of the learned dynamical models. To overcome this technical hurdle, we propose a new method that learns nonlinear dynamics through a Bayesian inference of characterizing model parameters. This method leverages a Gaussian process representation of states, and constructs a likelihood function using the correlation between state data and their derivatives, yet prevents explicit evaluations of time derivatives. Through a Bayesian scheme, a probabilistic estimate of the model parameters is given by the posterior distribution, and thus a quantification is facilitated for uncertainties from noisy state data and the learning process. Specifically, we will discuss the applicability of the proposed method to several typical scenarios for dynamical systems: identification and estimation with an affine parametrization, nonlinear parametric approximation without prior knowledge, and general parameter estimation for a given dynamical system.
@@ 1 Introduction
Ideally, physical consistency should be preserved in the learning process through a proper system parametrization, so that non-physical behaviors in the data are filtered out and reliability in long-term predictions is achieved. One popular choice is the sparse identification of nonlinear dynamics (SINDy) [7], which defines $f_{i}(\cdot;{\bf\it\theta}_{i})$ as a linear combination of candidate functions and uses sparse regression to determine active terms among them. The effectiveness of identification in this method relies on the inclusiveness of candidates and the assumption of sparsity. Relevantly, data-driven operator inference [28, 30, 19] focuses on reduced-order dynamics learning, and attempts to recover a low-dimensional system (1) with a polynomial structure $f_{i}(\cdot;{\bf\it\theta}_{i})$ inspired by known governing equations. Formulated as (Tikhonov-regularized) least squares [25], such a non-intrusive method seeks to find estimators of ${\bf\it\theta}_{i}$ that best match the training data in a minimum-residual sense, and hence avoids entailing access to source code. This is especially advantageous for applications with readily-executed solvers that cannot be easily modified. Alternative to pinpointing the exact structure of a dynamical system, $f_{i}(\cdot;{\bf\it\theta}_{i})$ can be considered as an approximation of the right-hand side of the differential equation with parameters ${\bf\it\theta}_{i}$. For example, neural ordinary differential equations [10] adopt a multi-layer perceptron network as $f_{i}(\cdot;{\bf\it\theta}_{i})$, while the Runge-Kutta neural network method [42] constructs parametric surrogate models for the discrete-time integrator.
@@
For many of these dynamics learning techniques, however, predictive performance can be compromised when training data are scarce and/or corrupted by noise. In particular with the ‘equation-fitting’ strategy, i.e., minimum-residual of the continuous- or discrete-time formulation of (1), there typically involve time derivative approximations by numerical differentiation (e.g., finite differences). The data noise and/or sparsity may lead to a poor estimation of time derivatives and then curtail the accuracy of parameter inference. Efforts have been made to avoid derivative evaluations by considering the misfit in state data (‘trajectory-fitting’) directly instead of minimizing residuals, such as incorporating time integration into loss functions [10] or as constraints [37]. However, these settings typically demand more sophisticated numerical treatments, which are not always convenient, flexible, or effective for complex dynamical behaviors. Meanwhile, particularly for the cases with imperfect datasets, it is crucial to facilitate the learning process with uncertainty quantification and enable a model validation in terms of accuracy and robustness. A straightforward strategy is ensemble learning that leverages bootstrap sampling of time-series data to aggregate statistical estimates from multiple models [15], but the many times of evaluations may escalate computational costs. An efficient and rigorous alternative is Bayesian inference [6] for system parameters ${\bf\it\theta}_{i}$, deemed capable of taking data and modeling uncertainties into systematical consideration [35, 41, 39, 21, 19]. In this work, we aim to address the aforementioned two challenges – robustness with scarce/noisy data and capability of uncertainty quantification – through a novel Bayesian inference method for learning dynamical systems (1). The primary novelty lies in integrating Gaussian process modeling into Bayesian inference to smooth and/or interpolate imperfect data, accounting for the approximation and uncertainty estimation of time derivatives, and incorporating differential equation constraints into the inference process.
@@
Gaussian process emulation is a probabilistic approach to supervised learning [32], and its variations have been applied to a wide range of computational tasks in science and engineering, see [4, 24, 14, 23, 9, 13, 40, 5, 12] for a few examples. More relevant to this work is the fact that Gaussian processes are also capable of embedding differential equation constraints into kernel structures. Specifically, as Gaussianity can be preserved through linear operations, the latent force models [43] for Gaussian processes have paved the way for describing the correlations between solution states/fields and their latent ‘forces’ through linear differential operators. This idea has been used for solving forward and inverse problems governed by linear partial differential equations [31, 29], and extended to nonlinear equations [11, 26] with error analysis [1], as well as to nonlinear operator learning [2]. Relevant techniques have been explored in various contexts in scientific computing, such as stochastic partial differential equations [33], statistical finite element analysis [17], port-Hamiltonian systems [3], and conservation laws [20].
@@
In this work, a vector-valued Gaussian process is constructed to represent the statistical correlation between the states and their time derivatives. Combined with the system parametrization $f_{i}(\cdot;{\bf\it\theta}_{i})$, this correlation is leveraged to formulate a novel likelihood function constrained by the differential equations (1), in addition to a prior distribution of the parameters ${\bf\it\theta}_{i}$ needed for the Bayesian inference. Thereafter, a probabilistic estimate of ${\bf\it\theta}_{i}$ is given by the posterior distribution through the Bayes’ rule. By sampling over this distribution, uncertainties are consequently propagated into the predictions produced by the learned dynamical system, enabling a predictive uncertainty quantification. Importantly, the Gaussian process representation also provides a natural smoother by kernel regression, which counters the unsatisfactory accuracy in time derivatives caused by noisy/scarce state data. We will further explain the proposed method in two scenarios: one with a given affine structure of the dynamical system, and the other for a general nonlinear approximation of the system without prior knowledge.
@@ 2 Dynamics learning with Gaussian processes
$\pi_{\texttt{prior}}({\bf\it\theta}_{i})\propto\exp\left(-\frac{1}{2}\|{\bf\it\theta}_{i}\|_{{\bf\it\Lambda}_{i}}^{2}\right)\,,$ (2)
@@ 2.1 Likelihood definition with a Gaussian process
Here we assume that each component $x_{i}(t)$ ($1\leq i\leq N$) of the state vector ${\bf\it x}(t)$ follows a zero-mean Gaussian process over the time $t$:
@@
in which $\kappa_{i}$ is a given smooth kernel function representing the covariance over the Gaussian process, i.e., $\kappa_{i}(t,t^{\prime})=\mathbb{C}ov[x_{i}(t),x_{i}(t^{\prime})]$. Particularly, we use the squared exponential kernel as the covariance function in this work. Thus, the time-derivative of $x_{i}(t)$ also follows a Gaussian process, and the correlation between the state $x_{i}$ and its time-derivative $\dot{x}_{i}$ is given by
@@
For notation, we collect the observables of $x_{i}(t)$ at time-instances $\mathcal{T}$ in a size-$K$ random vector ${\bf\it U}_{i}$, while their observed values correspondingly in $\mathbf{u}_{i}\in\mathbb{R}^{K}$, let ${\bf\it D}_{i}$ denote the size-$K$ random vector of $x_{i}$’s time derivatives over $\mathcal{T}$, and let $\mathbf{Z}=[\mathbf{u}_{1}~{}\cdots~{}\mathbf{u}_{N}]^{\mathsf{T}}=[{\bf\it x}(t_{0})~{}\cdots~{}{\bf\it x}(t_{K-1})]\in\mathbb{R}^{N\times K}$ collect the full-state data of ${\bf\it x}(t)$ over $\mathcal{T}$. As a Gaussian process assumption on $x_{i}(t)$ has been adopted, ${\bf\it D}_{i}$ combined with ${\bf\it U}_{i}$ should follow a joint normal distribution written as
$\begin{split}\begin{bmatrix}\boldsymbol{D}_{i}\\ \boldsymbol{U}_{i}\end{bmatrix}&\sim\mathcal{N}\left(\begin{bmatrix}\mathbf{0}\\ \mathbf{0}\end{bmatrix},\begin{bmatrix}\mathbf{K}^{dd}_{i}&\mathbf{K}^{du}_{i}\\ \mathbf{K}^{ud}_{i}&\mathbf{K}^{uu}_{i}\end{bmatrix}\right)\,,\quad\text{with}\\ \mathbf{K}_{i}^{dd}&=\partial_{t}\partial_{t^{\prime}}\kappa_{i}(\mathcal{T},\mathcal{T})+\chi^{d}_{i}\mathbf{I}_{K}\in\mathbb{R}^{K\times K}\,,\\ \mathbf{K}_{i}^{uu}&=\kappa_{i}(\mathcal{T},\mathcal{T})+\chi^{u}_{i}\mathbf{I}_{K}\in\mathbb{R}^{K\times K}\,,\\ \mathbf{K}_{i}^{du}&=\partial_{t}\kappa_{i}(\mathcal{T},\mathcal{T})=(\mathbf{K}_{i}^{ud})^{\mathsf{T}}\in\mathbb{R}^{K\times K}\,.\end{split}$ (5)
Considering the possibility that the data $\mathbf{Z}$ may be corrupted by noise, we have added white noise terms with variance values $\chi^{u}_{i}>0$ and $\chi^{d}_{i}>0$ to ${\bf\it U}_{i}$ and ${\bf\it D}_{i}$, respectively. Furthermore, we define an additional random vector ${\bf\it\delta}_{i}={\bf\it D}_{i}-f_{i}(\mathbf{Z};{\bf\it\theta}_{i})$ that represents the residuals over $\mathcal{T}$ with given state data $\mathbf{Z}$ and a fixed ${\bf\it\theta}_{i}$.
In this work, the core idea is to construct a likelihood function not only on the direct observations, but also on the alignment between ${\bf\it D}_{i}$ and $f_{i}(\mathbf{Z};{\bf\it\theta}_{i})$, i.e., ${\bf\it\delta}_{i}=\mathbf{0}$, the satisfaction of the dynamical system formulation (1). Such a likelihood function can thus be formulated from (5) as follows:
$\begin{split}\pi_{\texttt{like}}({\bf\it\theta}_{i})&=p\left({\bf\it\delta}_{i}=\mathbf{0},{\bf\it U}_{i}=\mathbf{u}_{i}\mid\mathbf{Z},{\bf\it\theta}_{i}\right)\\ &\propto\exp\left(-\frac{1}{2}\begin{bmatrix}f_{i}(\mathbf{Z};{\bf\it\theta}_{i})\\ \mathbf{u}_{i}\end{bmatrix}^{\mathsf{T}}\begin{bmatrix}\mathbf{K}^{dd}_{i}&\mathbf{K}^{du}_{i}\\ \mathbf{K}^{ud}_{i}&\mathbf{K}^{uu}_{i}\end{bmatrix}^{\text{-1}}\begin{bmatrix}f_{i}(\mathbf{Z};{\bf\it\theta}_{i})\\ \mathbf{u}_{i}\end{bmatrix}\right)\,.\end{split}$ (6)
Therefore, in addition to the data-driven nature, such a likelihood definition guarantees a physics-based interpretation.
Remark 1: In this work, we consider both ${\bf\it U}_{i}$ and ${\bf\it D}_{i}$ over the whole data coverage $\mathcal{T}$, but this is not a requirement, i.e., ${\bf\it U}_{i}$ and ${\bf\it D}_{i}$ can correspond to two different subsets of time instances in $\mathcal{T}$.
@@ 2.2 Posterior distribution for Bayesian inference
Note that the inverse of the kernel matrix in the likelihood (6) can be rewritten as
@@
By substituting this into the $\pi_{\texttt{like}}$ expression (6) that is used in the inference (7), the posterior distribution is therefore written explicitly as
$\begin{split}&~{}~{}\pi_{\texttt{post}}({\bf\it\theta}_{i})\propto\pi_{\texttt{like}}({\bf\it\theta}_{i})\pi_{\texttt{prior}}({\bf\it\theta}_{i})\\ \propto&~{}\exp\left(-\frac{1}{2}\left(f_{i}(\mathbf{Z};{\bf\it\theta}_{i})^{\mathsf{T}}\mathbf{R}_{i}^{dd}f_{i}(\mathbf{Z};{\bf\it\theta}_{i})+2f_{i}(\mathbf{Z};{\bf\it\theta}_{i})^{\mathsf{T}}\mathbf{R}_{i}^{du}\mathbf{u}_{i}+\|{\bf\it\theta}_{i}\|_{{\bf\it\Lambda}_{i}}^{2}\right)\right)\,.\end{split}$ (9)
Additionally, we define an estimate of the time derivatives over $\mathcal{T}$ (i.e., $\dot{x}_{i}(\mathcal{T})$) using a Gaussian process regression with the state observations $(\mathcal{T},\mathbf{u}_{i})$ at the same instances, given as
@@
These are the time derivatives extracted from the mean function $\kappa_{i}(t,\mathcal{T})(\mathbf{K}_{i}^{uu})^{-1}\mathbf{u}_{i}$ of the Gaussian process reconstruction. Considering the fact that $\mathbf{R}_{i}^{dd}\hat{\mathbf{d}}_{i}=-\mathbf{R}_{i}^{du}\mathbf{u}_{i}$, the posterior expression in (9) is hence rearranged into
$\pi_{\texttt{post}}({\bf\it\theta}_{i})\propto\exp\left(-\frac{1}{2}\left(\left\|f_{i}(\mathbf{Z};{\bf\it\theta}_{i})-\hat{\mathbf{d}}_{i}\right\|_{\mathbf{R}_{i}^{dd}}^{2}+\|{\bf\it\theta}_{i}\|_{{\bf\it\Lambda}_{i}}^{2}\right)\right)\,,$ (11)
@@
$({\bf\it\theta}_{i})_{\text{MAP}}=\arg\min_{{\bf\it\theta}_{i}\in\mathbb{R}^{p_{i}}}\left\{\left\|f_{i}(\mathbf{Z};{\bf\it\theta}_{i})-\hat{\mathbf{d}}_{i}\right\|_{\mathbf{R}_{i}^{dd}}^{2}+\|{\bf\it\theta}_{i}\|_{{\bf\it\Lambda}_{i}}^{2}\right\}\,.$ (12)
This is a regularized, generalized least squares estimator of ${\bf\it\theta}_{i}$ weighted by $\mathbf{R}_{i}^{dd}=(\mathbf{K}^{dd}_{i}-\mathbf{K}^{du}_{i}(\mathbf{K}^{uu}_{i})^{-1}\mathbf{K}^{ud}_{i})^{-1}$, the inverse of which is the posterior covariance matrix of $\dot{x}_{i}(\mathcal{T})$ through Gaussian process regression. Inside the first term of the loss function is the residual of the governing differential equation over $\mathcal{T}$ with time derivatives approximated by Gaussian process regression. Essentially, this confirms the interpretability of the proposed method. With the weight matrix $\mathbf{R}_{i}^{dd}$, the uncertainty introduced by the derivative approximation is included into the posterior estimation.
@@
$\pi_{\texttt{post}}({\bf\it\theta}_{i})\propto\exp\left(-\frac{1}{2}\left(\sum_{\texttt{all ICs}}\left\{\left\|f_{i}(\mathbf{Z};{\bf\it\theta}_{i})-\hat{\mathbf{d}}_{i}\right\|_{\mathbf{R}_{i}^{dd}}^{2}\right\}_{\texttt{each IC}}+\|{\bf\it\theta}_{i}\|_{{\bf\it\Lambda}_{i}}^{2}\right)\right)\,,$ (13)
where the loss terms $\|f_{i}(\mathbf{Z};{\bf\it\theta}_{i})-\hat{\mathbf{d}}_{i}\|_{\mathbf{R}_{i}^{dd}}^{2}$ are summed over all the initial conditions (ICs) included in the dataset. Note in this case that the trajectory with each initial condition is approximated by an individual GP, leading correspondingly to an individual $\mathbf{R}_{i}^{dd}$ as well. In particular, when data with a single initial condition are scarce, one can try to improve the parameter inference by adding data with other initial conditions. The estimated parameter values should work for the predictions for new initial conditions outside the coverage of training data.
@@
$\begin{split}&\pi_{\texttt{post}}({\bf\it\theta}_{m}\cup{\bf\it\theta}_{n})=p\left({\bf\it\theta}_{m}\cup{\bf\it\theta}_{n}\mid{\bf\it\delta}_{m}=\mathbf{0},~{}{\bf\it\delta}_{n}=\mathbf{0},~{}{\bf\it U}_{m}=\mathbf{u}_{m},~{}{\bf\it U}_{n}=\mathbf{u}_{n},~{}\mathbf{Z}\right)\\ \propto&\exp\left(-\frac{1}{2}\left(\left\|f_{m}(\mathbf{Z};{\bf\it\theta}_{m})-\hat{\mathbf{d}}_{m}\right\|_{\mathbf{R}_{m}^{dd}}^{2}+\left\|f_{n}(\mathbf{Z};{\bf\it\theta}_{n})-\hat{\mathbf{d}}_{n}\right\|_{\mathbf{R}_{n}^{dd}}^{2}+\|{\bf\it\theta}_{i}\|_{{\bf\it\Lambda}_{m}}^{2}+\|{\bf\it\theta}_{i}\|_{{\bf\it\Lambda}_{n}}^{2}\right)\right)\,.\end{split}$ (14)